# Emergent time and more from wavefunction collapse in general relativity

### v1

@@ II Wavefunction collapse from stochastic fluctuations of lapse and shift
In obtaining this expression, it is convenient to use the normal ordering in Eq. (4). In different normal orderings, $A_{0}$ is modified, but one also needs to include additional boundary terms due to the essentially non-self-adjoint nature of $\hat{\cal H}$ (see Appendix A for details). A gradual collapse of the initial state toward ${\mathbb{H}}_{0}$ is achieved through a stochastic sum of the lapse and shift functions as
$|\Psi\rangle$ $=$ $\int{\cal D}g{\cal D}\Pi~~|g^{(N)}\rangle\int\prod_{m=1}^{N}\left[D\epsilon^{(m)}{\cal P}^{(m)}(\epsilon^{(m)})\right]e^{iS}$$\times$ (6)
$\Psi_{0}(g^{(0)}),$
where ${\cal P}^{(m)}(\epsilon)$ represents the Gaussian distribution of the lapse and shift with zero mean and a non-zero second moment: $\int D\epsilon{\cal P}^{(m)}(\epsilon)\epsilon_{0}(x)=\int D\epsilon{\cal P}^{(m)}(\epsilon)\epsilon_{\mu}(x)=\int D\epsilon{\cal P}^{(m)}(\epsilon)\epsilon_{0}(x)\epsilon_{\mu}(y)=0$, $\int D\epsilon{\cal P}^{(m)}(\epsilon)\epsilon_{0}(x)\epsilon_{0}(y)=\frac{2\epsilon^{2}\xi^{(m)}(x)}{\sqrt{g^{(m)}}}\delta(x-y)$, and $\int D\epsilon{\cal P}^{(m)}(\epsilon)\epsilon_{\mu}(x)\epsilon_{\nu}(y)=\frac{\epsilon^{2}\xi^{{}^{\prime}(m)}(x)g^{(m)}_{\mu\nu}}{2\sqrt{g^{(m)}}}\delta(x-y)$. The second moment is proportional to $\epsilon^{2}$. $\xi^{(m)}(x)$ and $\xi^{{}^{\prime}(m)}(x)$ are functions that characterize the speed of collapse, which is generally position-dependent. In the large $N$ limit with fixed $L\equiv N\epsilon^{2}$, we obtain a $L$-dependent state given by
@@ IV Excitations
In the the adiabatic limit, the effective Hamiltonian is obtained to be (see Appendix B for details)
@@
$+is\int dx\frac{1}{\sqrt{\bar{g}(t)}}\left[C_{\Phi}(t)(\hat{\pi}^{\prime}_{\Phi})^{2}+C_{\gamma_{\perp}}(t)\bar{g}^{\mu\nu}(t)\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right],$
@@ V Physical consequences
$F_{k}(t)=e^{-\frac{A|k(t)|}{(d+2)\Lambda}e^{\sqrt{\frac{\Lambda}{d(d-1)}}t}\left(e^{-\sqrt{\frac{d\Lambda}{4(d-1)}}t_{0}}-e^{-\sqrt{\frac{d\Lambda}{4(d-1)}}t}\right)}$ (45)
@@ VI Summary and Discussion
Here, we conclude with a few future directions. Firstly, one can try to look for signatures of non-unitarity in the scattering amplitudes. For example, the scattering amplitudes of the tensor graviton and the particles in the standard model will be affected by the short-lived, unstable particles. At low energies, the effect is expected to be small due to the weak interaction; however, there should be observable effects at sufficiently high energies. Secondly, it is of interest to extend the present study to configurations with strong spatial inhomogeneity. While the principle remains the same, the inhomogeneity may affect the progression of wavefunction collapse in a space-dependent manner, especially near compact objects such as black holes.
@@ Appendix B Effective Hamiltonian
$+is\int dxe^{-d\bar{\alpha}}\left\{\frac{1}{4}(\hat{\pi}_{\Phi})^{2}+\bar{g}^{\mu\nu}\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right\}.$ (55)
@@
where $f_{k}=\frac{e^{d\bar{\alpha}}|k(l;L)|}{2}$ and $f_{\chi,k}=\sqrt{is}\sqrt{\frac{(d-1)(d-2)|{\cal N}|}{u\Lambda\left(1+\frac{|k(l;L)|^{2}}{4d^{2}u\Lambda}\right)}}e^{d\bar{\alpha}}|k(l;L)|$. In $f_{\chi,k}$, the sign ambiguity in $\sqrt{is}$ should be resolved such that the vacuum wavefunction $\Psi_{vac}(\chi)\sim e^{-\frac{1}{2}\sum_{k}f_{\chi,k}\chi_{k}\chi_{-k}}$ annihilated by $\hat{b}_{k}$ is normalizable. This forces the real part of $f_{\chi,k}$ to be positive and $\sqrt{is}=e^{is\pi/4}=\frac{1}{\sqrt{2}}(1+is)$. While $\hat{a}^{(j)\dagger}_{k}$ is the adjoint of $\hat{a}^{(j)}_{k}$, $\hat{b}_{k}^{\dagger}$ is not the adjoint of $\hat{b}_{k}$ because $f_{\chi,k}$ is not real. This is expected because the Hamiltonian that governs the scalar is not a multiple of a Hermitian operator.
@@
$+is\int dx\frac{1}{\sqrt{\bar{g}(t)}}\left[C_{\Phi}(t)(\hat{\pi}^{\prime}_{\Phi})^{2}+C_{\gamma_{\perp}}(t)\bar{g}^{\mu\nu}(t)\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right],$

### v2

@@ II Wavefunction collapse from stochastic fluctuations of lapse and shift
In obtaining this expression, it is convenient to use the normal ordering in Eq. (4). In different normal orderings, $A_{0}$ is modified, but one also needs to include additional boundary terms due to the essentially non-self-adjoint nature of $\hat{\cal H}$ (see Appendix A for details). To describe a gradual collapse of the initial state toward ${\mathbb{H}}_{0}$, we write the infinitesimal gauge parameters as $\epsilon_{0}^{(m)}(x)=\varepsilon n_{0}^{(m)}(x)$ and $\epsilon_{\mu}^{(m)}(x)=\varepsilon n_{\mu}^{(m)}(x)$. Here, $\varepsilon$ is an infinitesimal parameter, and $n_{0}^{(m)}(x)$ and $n_{\mu}^{(m)}(x)$ are, respectively, the lapse and shift functions. The collapse arises from fluctuations of the lapse and shift functions,
$|\Psi\rangle$ $=$ $\int{\cal D}g{\cal D}\Pi|g^{(N)}\rangle\left[\int{\cal D}n\rho(n)e^{iS}\right]\Psi_{0}(g^{(0)}),$ (6)
where $\int{\cal D}n=\int\left[\prod_{m=1}^{N}Dn^{(m)}\right]$ averages over the lapse and shift functions over the distribution $\rho(n)$. To realize an unbiased collapse that does not have a preferred direction, we demand that $\rho(n)$ is even in $n_{0}$ and $n_{\mu}$. We also assume that the lapse and shift have zero correlation length in space. We write the second moments of the distribution as $\int{\cal D}n\rho(n)n^{(m)}_{0}(x)n^{(m^{\prime})}_{0}(y)=\frac{2\xi^{(m)}(x)}{\varepsilon\sqrt{g^{(m)}}}\delta(x-y)\delta_{m,m^{\prime}}$, $\int{\cal D}n\rho(n)n^{(m)}_{\mu}(x)n^{(m^{\prime})}_{\nu}(y)=\frac{\xi^{{}^{\prime}(m)}(x)g^{(m)}_{\mu\nu}}{2\varepsilon\sqrt{g^{(m)}}}\delta(x-y)\delta_{m,m^{\prime}}$ and $\int{\cal D}n\rho(n)n^{(m)}_{0}(x)n^{(m^{\prime})}_{\nu}(y)=0$. Here, $\xi^{(m)}(x)$ and $\xi^{{}^{\prime}(m)}(x)$ are functions that characterize the speed of collapse, which is generally position-dependent. The average over the lapse and shift gives
$\int{\cal D}n~\rho(n)~e^{-i\varepsilon\sum_{m}\int dx\left[n_{0}^{(m)}{\cal H}^{(m)}+n_{\mu}^{(m)}{\cal P}^{(m)\mu}\right]}=$
$\exp\Bigg\{-\sum_{m}\int dx\frac{\varepsilon}{\sqrt{g^{(m)}}}\Bigg[\xi^{(m)}({\cal H}^{(m)})^{2}+\xi^{{}^{\prime}(m)}({\cal P}^{(m)})^{2}$
$~~~~~~~~~~+{\cal O}\left(\frac{({\cal H}^{(m)})^{4}}{g^{(m)}}\right)+{\cal O}\left(\frac{({\cal P}^{(m)})^{4}}{g^{(m)}}\right)\Bigg]\Bigg\},$ (7)
[…]
@@ IV Excitations
In the adiabatic limit, the effective Hamiltonian is obtained to be (see Appendix B for details)
@@
$+is\int dx\frac{1}{\sqrt{\bar{g}(t)}}\left[C_{\Phi}(t)(\hat{\pi}^{\prime}_{\Phi})^{2}+C_{\gamma_{\perp}}(t)\bar{g}^{\mu\nu}(t)\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right].$
@@ V Physical consequences
$F_{k}(t)=e^{-\frac{A|k(t)|}{(d+2)\Lambda}e^{\sqrt{\frac{\Lambda}{d(d-1)}}t}\left(e^{-\sqrt{\frac{d\Lambda}{4(d-1)}}t_{0}}-e^{-\sqrt{\frac{d\Lambda}{4(d-1)}}t}\right)}.$ (46)
@@
Here, the large $d$ limit is considered to control our calculations. In $d=3$, the $1/d$ corrections can be potentially significant. In Appendix C, we identify the sources of such $1/d$ corrections.
@@ VI Summary and Discussion
Here, we conclude with a few future directions. Firstly, one can try to look for signatures of non-unitarity in the scattering amplitudes. For example, the scattering amplitudes of the tensor graviton and the particles in the standard model will be affected by the short-lived, unstable particles. At low energies, the effect is expected to be small due to the weak interaction; however, there should be observable effects at sufficiently high energies. Secondly, it is of interest to extend the present study to configurations with strong spatial inhomogeneity. While the principle remains the same, the inhomogeneity may affect the progression of wavefunction collapse in a space-dependent manner, especially near compact objects such as black holes. Finally, one can consider more general cases in which the cosmological constant is not positive and the space has non-zero curvature. Around a homogeneous and isotropic space, the curvature scalar can be written as $R=Ke^{-2\alpha}+\tilde{R}$, where the first term originates from the curvature of the background space, and $\tilde{R}$ is the contribution of fluctuations. Then, the Hamiltonian density is written as
${\cal H}=u\Bigl[\tilde{\Lambda}(\alpha)+e^{-2d\alpha}G_{\mu\nu\rho\sigma}\tilde{\Pi}^{\mu\nu}\tilde{\Pi}^{\rho\sigma}-\tilde{R}\Bigr]-e^{-2d\alpha}\Pi^{2},$
where $\tilde{\Lambda}(\alpha)=\Lambda-Ke^{-2\alpha}$ is the effective vacuum energy. For the saddle-point without excitation, we can set $G_{\mu\nu\rho\sigma}\tilde{\Pi}^{\mu\nu}\tilde{\Pi}^{\rho\sigma}-\tilde{R}=0$. Even for $\Lambda<0$, $\tilde{\Lambda}(\alpha)$ needs to be non-negative; if $\tilde{\Lambda}(\alpha)<0$, there is no configuration that satisfies ${\cal H}=0$ and the wavefunction collapse results in $\Psi=0$ in the late time limit. If an initial state satisfies $\tilde{\Lambda}(\alpha)>0$ with a sufficiently negative $K$, the evolution generated by the wavefunction collapse is similar to the case with a positive cosmological constant. In particular, $\alpha$ still increases under the wavefunction collapse, and the direction of time evolution is determined by the initial state in the same way. However, the dependence of $\bar{\alpha}$ on $L$ will be modifiedLee (2025) because $\tilde{\Lambda}(\alpha)$ decreases with increasing $\alpha$. More importantly, a qualitatively new behavior is expected to arise as the universe approaches the critical size $\alpha_{c}$ at which $\tilde{\Lambda}(\alpha_{c})=0$. Near $\alpha_{c}$, the higher order terms in $\partial_{l}\alpha/\tilde{\Lambda}(\alpha)$ play important roles in determining the dynamics of $\alpha$ in Eq. (13). The same issue arises in flat space with $\Lambda=0$. It would be of great interest to understand how the late time dynamics differ in the spaces with $\Lambda\leq 0$ compared to the one with $\Lambda>0$.
@@ Appendix B Effective Hamiltonian
$+is\int dxe^{-d\bar{\alpha}}\left\{\frac{1}{4}(\hat{\pi}_{\Phi})^{2}+\bar{g}^{\mu\nu}\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right\},$ (57)
@@
where $f_{k}=\frac{e^{d\bar{\alpha}}|k(l;L)|}{2}$ and $f_{\chi,k}=\sqrt{is}\sqrt{\frac{(d-1)(d-2)|{\cal N}|}{u\Lambda\left(1+\frac{|k(l;L)|^{2}}{4d^{2}u\Lambda}\right)}}e^{d\bar{\alpha}}|k(l;L)|$. In $f_{\chi,k}$, the sign ambiguity in $\sqrt{is}$ should be resolved such that the vacuum wavefunction $\Psi_{vac}(\chi)\sim e^{-\frac{1}{2}\sum_{k}f_{\chi,k}\chi_{k}\chi_{-k}}$ annihilated by $\hat{b}_{k}$ is normalizable. This forces the real part of $f_{\chi,k}$ to be positive and $\sqrt{is}=e^{is\pi/4}=\frac{1}{\sqrt{2}}(1+is)$. While $\hat{a}^{(j)\dagger}_{k}$ is the adjoint of $\hat{a}^{(j)}_{k}$, $\hat{b}_{k}^{\dagger}$ is not the adjoint of $\hat{b}_{k}$ because $f_{\chi,k}$ is not real. This is because the Hamiltonian that governs the scalar is not a multiple of a Hermitian operator.
@@
$+is\int dx\frac{1}{\sqrt{\bar{g}(t)}}\left[C_{\Phi}(t)(\hat{\pi}^{\prime}_{\Phi})^{2}+C_{\gamma_{\perp}}(t)\bar{g}^{\mu\nu}(t)\hat{\pi}_{\gamma_{\perp}\mu}\hat{\pi}_{\gamma_{\perp}\nu}\right].$
@@
Appendix C $1/d$ corrections
In this appendix, we revisit the approximations employed in the large $d$ limit and discuss how $1/d$ corrections arise at finite $d$.
  • Gradient expansion of the saddle point;
    In Eq. (13), the saddle-point solution for $\Pi$ only keeps the term to the linear order in $\partial_{l}\bar{\alpha}/P^{3}$. This approximation is justified if
    $\partial_{l}\bar{\alpha}\ll(u\Lambda)^{3/2},$ (64)
    where u=d(d1)u=d(d-1). According to Eq. (25),
    $\partial_{l}\bar{\alpha}\leq\frac{\left(\sqrt{1+4Aud\Lambda Le^{-d\alpha_{0}}}-1\right)}{dL},$ (65)
    and Eq. (64) is guaranteed to be satisfied in the large $d$ limit. For a finite $d$ with $\sqrt{u\Lambda}<2Ae^{-d\alpha_{0}}$, Eq. (64) is violated at small $L$. In that case, the higher order terms in $\partial_{l}\alpha/P^{3}$ become important for the ‘early’ universe with $L\leq\frac{2e^{-d\alpha_{0}}\left(2A-e^{d\alpha_{0}}\sqrt{\Lambda u}\right)}{d\Lambda^{2}u^{2}}$.
  • Adiabatic approximation;
    Eq. (42) is obtained based on the adiabatic approximation, which is valid for the modes whose frequencies in $l$ are much greater than that rate at which $\bar{\alpha}$ changes,
    $\partial_{l}\bar{\alpha}\ll\omega_{k}(l;L),\omega^{\chi}_{k}(l;L),$ (66)
    where $\omega_{k}(l;L)$ and $\omega^{\chi}_{k}(l;L)$ are given in Eq. (59). Eq. (66) is satisfied for the modes with $|k(l;L)|\gg\sqrt{\Lambda/u},~(\partial\bar{\alpha})^{1/2}\Lambda^{-1/4}u^{-5/4}$, where $|k(l;L)|=\sqrt{\bar{g}^{\mu\nu}(l;L)k_{\mu}k_{\nu}}$. In the large $d$ limit, the adiabatic approximation is valid for all non-zero $k$ both for the tensor and scalar gravitons. For a finite $d$, one should include non-adiabatic processes in which the expanding geometry creates pairs of particles at low momenta below crossover momentum scales. For the tensor graviton, the crossover momentum is fixed to be $\sqrt{\Lambda/u}$. For the scalar graviton, the crossover momentum increases with decreasing $l$ as the universe expands faster at earlier times. Therefore, the non-adiabatic corrections are present over a larger range of momenta at smaller $l$. However, the crossover momentum is still bounded from above because Eq. (65) is bounded.
  • Quadratic approximation;
    Eq. (42) has been derived from the quadratic action. In general, the cubic and higher-order interaction vertices in Eq. (11) renormalize the effective action, including the kinetic energy. Since the vector graviton is non-propagating at long distance scales, we focus on the scalar and tensor gravitons. The Lagrangian density that includes the cubic and quartic vertices for the propagating modes can be written as
    ${\cal L}^{\prime}$ $\sim$ $\delta\pi\partial_{l}\varphi+\pi_{\perp}^{\mu\nu}\partial_{l}h_{\mu\nu}-{\cal N}\left[{\bf H}_{\perp}+D_{4}{\bf H}_{\perp}^{2}\right]$
    $+i\left[C_{2}\delta\pi^{2}+C_{3}\delta\pi^{3}+C_{4}\delta\pi^{4}\right],$
    ${\bf H}_{\perp}$ $\sim$ $\pi_{\perp,\nu}^{\mu}\pi_{\perp,\mu}^{\nu}-\frac{1}{4}h_{\nu}^{\mu}\bar{\nabla}^{2}h_{\mu}^{\nu}-B_{2}\left|\bar{\nabla}\varphi\right|^{2},$ (67)
    where $\delta\pi=e^{-d\bar{\alpha}}\delta\Pi$, $\pi_{\perp}^{\mu\nu}=e^{-d\bar{\alpha}}\Pi_{\perp}^{\mu\nu}$, $C_{2}\sim u\Lambda$, $C_{3}\sim\sqrt{u\Lambda}$, $C_{4}\sim 1$, $B_{2}\sim(d-1)(d-2)$, $D_{4}\sim\Lambda^{-1}$ and $\cal N$} is defined in Eq. (33). The non-quadratic terms inside $\bf H$$_{\perp}$ are dropped because they are suppressed below the Planck energy scale $l_{p}^{-1}$, which is set to be $1$ here. At low energies, the interactions from ${\bf H}_{\perp}^{2}$ are more important than those.
    In the adiabatic limit, the scaling dimensions of the fields are given by $\varphi]=[h^{\mu}_{\nu}]=(d-1)/2$ and $\delta\pi]=[\pi^{\mu}_{\perp,\nu}]=(d+1)/2$. Under the scale transformation that leaves the quadratic terms invariant, $C_{3},C_{4},D_{4}$ carry negative scaling dimensions, and the quantum corrections from them become negligible below crossover energy scales set by the dimensionful coefficients. The self-energy for the scalar mode generated from the non-quadratic terms of $\delta\pi$ becomes negligible compared with the bare kinetic energy for $|k(l;L)|\ll\left[u^{\frac{3}{4}}\Lambda^{\frac{7}{4}}\partial_{l}\bar{\alpha}^{-\frac{1}{2}}\right]^{\frac{1}{d+1}}$. Here, the factor of $\partial_{l}\bar{\alpha}^{-\frac{1}{2}}$ arises due to the fact that the momentum scale is related to the frequency scale through $k\sim(\partial_{l}\bar{\alpha})^{-\frac{1}{2}}\omega$ in Eq. (59) for the scalar mode. On the other hand, the self-energy from ${\bf H}_{\perp}^{2}$ is suppressed for $|k(l;L)|\ll\Lambda^{\frac{1}{d+1}}$. In the large $d$ limit, the crossover energy scale becomes large and approaches the Planck scale. For finite $d$ and small $\Lambda$, however, the crossover energy scales become significantly lower than the Planck scale; for $d=3$ and $\Lambda\sim 10^{-120}$, the self-energy of the graviton becomes important above $10^{-30}$ in the Planck unit, which becomes $\sim 10^{-2}eV$ for $l_{p}^{-1}\sim 10^{28}eV$. This implies that the gravitons become incoherent above the crossover energy scale due to interactions.